A force of 15 N is applied parallel to a lever arm of length 10 m. What is the resulting torque?

  

0 Nm

5 Nm

75 Nm

150 Nm
Answer #1

We know that Torque can be calculated as  

  \tau = \widetilde{r} \times \widetilde{F}

\tau = rF \sin\theta

Here the Force is applied parallel to the Lever Arm i.e \theta= 0\degree

Therefore the torque

\tau=\boldsymbol{\textbf{0}} Nm

Hope this Helps